# Open dynamical systems with a moving hole

### v1

@@ 5. Applications in badly approximable numbers, non-recurrence points and finiteness property
Suppose that $m\in\mathbb{N}_{\geq 10}$. For any $\omega=\omega^{0}\omega^{1}\ldots,\tau=\tau^{0}\tau^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ we have
@@
(5.1) $\mathbf{d}^{(j)}\notin\bigcup_{i=0}^{q-1}\left\{\omega^{jq+i}_{q+1-i}\omega^{jq+i}_{q+2-i}\ldots\omega^{jq+i}_{q+\ell-i},\leavevmode\nobreak\ \tau^{jq+i}_{q+1-i}\tau^{jq+i}_{q+2-i}\ldots\tau^{jq+i}_{q+\ell-i}\right\}.$
@@
$K_{n}:=\left\{x\in[0,1):T_{b}^{k}x\notin B_{k}\quad\forall k\in\mathbb{N}_{\geq n}\right\}.$
@@
$\widetilde{K}_{n}:=T_{b}^{n}(K_{n})=\left\{y\in[0,1):T_{b}^{k}y\notin B_{k+n}\quad\forall k\in\mathbb{N}_{0}\right\}.$
@@
Write $r_{n}=\mathrm{diam}\leavevmode\nobreak\ (B_{n})$ for $n\in\mathbb{N}_{0}$. Suppose that $r_{n}\to 0$ as $n\to\infty$. Then for any $m\in\mathbb{N}_{\geq 10}$ we can find $n_{0}\in\mathbb{N}$ such that $r_{n_{0}+k}<b^{-m}$ for all $k\in\mathbb{N}_{0}$. Take $n\geq n_{0}$. Then for any $k\in\mathbb{N}_{0}$ we can find two words $\mathbf{c}_{k},\mathbf{d}_{k}\in D_{b}^{m}$ such that $B_{k+n}\subset I_{\mathbf{c}_{k}}\cup I_{\mathbf{d}_{k}}$. It follows that
(5.3) $\widetilde{K}^{\omega,\tau}:=\left\{x\in[0,1):T_{b}^{k}x\not\in I_{\mathbf{c}_{k}}\cup I_{\mathbf{d}_{k}}\quad\forall k\in\mathbb{N}_{0}\right\}\subset\widetilde{K}_{n},$
where $\omega=\mathbf{c}_{0}\mathbf{c}_{1}\mathbf{c}_{2}\ldots,\tau=\mathbf{d}_{0}\mathbf{d}_{1}\mathbf{d}_{2}\ldots\in\left(D_{b}^{m}\right)^{\mathbb{N}}$. Note that $\widetilde{K}^{\omega,\tau}$ and $K^{\omega,\tau}$ differ up to a countable set. By (5.3) and Lemma 5.1, we obtain that
@@
Next we assume that $\lim_{n\to\infty}r_{n}=r>0$. Choose $m\in\mathbb{N}$ such that $2b^{-m}<r$. Then there exists $n_{1}\in\mathbb{N}$ such that $r_{n_{1}+k}>2b^{-m}$ for all $k\in\mathbb{N}_{0}$. Take $n\geq n_{1}$. For each $k\in\mathbb{N}_{0}$ we can choose $\mathbf{d}_{k}\in D_{b}^{m}$ such that $I_{\mathbf{d}_{k}}\subset B_{k+n}$. Then we have
(5.4) $\widetilde{K}_{n}\subset\left\{x\in[0,1):T_{b}^{k}x\not\in\mathrm{int}\big{(}I_{\mathbf{d}_{k}}\big{)}\quad\forall k\in\mathbb{N}_{0}\right\}=:\widetilde{K}^{\omega},$
where $\omega=\mathbf{d}_{0}\mathbf{d}_{1}\mathbf{d}_{2}\ldots\in(D_{b}^{m})^{\mathbb{N}}$. Note that the sets $\widetilde{K}^{\omega}$ and $K^{\omega}$ differ up to a countable set. So, by Theorem 1.6 we obtain
@@
$E_{n}(\phi):=\left\{x\in[0,1):|T_{b}^{k}x-x|\geq\phi(k)\quad\forall k\geq n\right\}.$
@@
(5.5) $\dim_{H}E(\phi)=\lim_{n\to\infty}\dim_{H}E_{n}(\phi).$
First we prove the necessity. Suppose on the contrary that $\lim_{n\to\infty}\phi(n)>0$. Note that $\phi(n)>0$ for all $n\in\mathbb{N}$. Then $r:=\displaystyle\inf_{n\geq 1}\phi(n)>0$. Choose $m\in\mathbb{N}_{\geq 10}$ such that $b^{-m}<r$. Then for any word $\mathbf{d}=d_{1}d_{2}\ldots d_{m}\in D_{b}^{m}$, we have for $n\in\mathbb{N}$,
$E_{n}(\phi)\cap I_{\mathbf{d}}\subset\left\{x\in[0,1):T_{b}^{k}x\notin I_{\mathbf{d}}\quad\forall k\geq n\right\}\subset T_{b}^{-n}\widetilde{K}^{\omega}\quad\textrm{with}\quad\omega=\mathbf{d}^{\infty}\in\left(D_{b}^{m}\right)^{\mathbb{N}}.$
@@
Next we prove the sufficiency. Suppose $\phi(n)\to 0$ as $n\to\infty$. For any $m\in\mathbb{N}_{\geq 10}$ we can find $n_{0}\in\mathbb{N}$ such that $\phi(n)<b^{-m}$ for all $n\geq n_{0}$. For $n\geq\max\{m,n_{0}\}$ we have
$E_{n}(\phi)\cap[0,b^{-m}]\supset\left\{x\in[0,b^{-m}]:T_{b}^{k}x\not\in[0,2b^{-m})\quad\forall k\geq n\right\}\supset\widetilde{E}_{n},$
where
By Lemma 5.1 we obtain that
$\dim_{H}E_{n}(\phi)\geq\dim_{H}E_{n}(\phi)\cap[0,b^{-m}]\geq\dim_{H}\widetilde{E}_{n}\geq\dim_{H}(T^{n}\widetilde{E}_{n})\geq 1-\frac{2}{m-2}\Big{(}\frac{\log m}{\log b}+1\Big{)}.$
It follows from (5.5) that
Letting $m\to\infty$ we conclude that $\dim_{H}E\left(\phi\right)=1$ as desired. ∎
@@
On the other hand, suppose that $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots=(\omega^{0}\omega^{1}\ldots\omega^{n-1})^{\infty}\in\left(D_{b}^{m}\right)^{\mathbb{N}}$ is a periodic and progressively overlapping sequence. By Lemma 2.3 and Proposition 3.5, we have

### v2

@@ 5. Regularity of $K^{\omega}$
For $p,q\in\mathbb{N}$, the matrix $A^{p}B^{q}$ is a primitive nonnegative matrix.
Proof.
In order to describe the matrix $B^{k}$ for $k\in\mathbb{N}$, we define the sequence $\{v_{n}\}$ as follows:
$\begin{cases}v_{-2m}=\cdots=v_{-m-1}=0,\;v_{-m}=-1,\;v_{-m+1}=\cdots=v_{-1}=0,\\ v_{n}=bv_{n-1}-v_{n-m}\;\text{for}\;n\in\mathbb{N}_{0}.\end{cases}$
Then we have $v_{n}=b^{n}$ for $0\leq n\leq m-1$. One can check that
  1. (i)
    the sequence $\{v_{n}\}_{n=0}^{\infty}$ is positive and is strictly increasing, and
    $v_{n}>\max\{v_{n-1},v_{n-2},\ldots,v_{n-m+1}\}\quad\forall n\geq 0,$
  2. (ii)
    the sequence $\{v_{n}\}_{n=-m+1}^{\infty}$ is nonnegative, and
    $v_{n}=bv_{n-1}-v_{n-m}\geq\max\{v_{n-1},v_{n-2},\ldots,v_{n-m+1}\}\quad\forall n\geq-m+2.$
Write $B^{k}=(b_{i,j}^{(k)})_{1\leq i,j\leq m}$ for $k\in\mathbb{N}$. We claim that for $k\in\mathbb{N}$,
(5.2) $B^{k}=\left(\begin{array}[]{ccccc}v_{k}&v_{k-1}&\cdots&v_{k-m+2}&v_{k-m+1}\\ -v_{k-m+1}&-v_{k-m}&\cdots&-v_{k-2m+3}&-v_{k-2m+2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ -v_{k-2}&-v_{k-3}&\cdots&-v_{k-m}&-v_{k-m-1}\\ -v_{k-1}&-v_{k-2}&\cdots&-v_{k-m+1}&-v_{k-m}\end{array}\right)_{m\times m}.$
That is, $b_{1,j}^{(k)}=v_{k-j+1}$ for $1\leq j\leq m$, and $b_{i,j}^{(k)}=-v_{k+i-j-m}$ for $1<i\leq m,1\leq j\leq m$.
We will prove the claim (5.2) by induction on $k$. First, by (3.10) the claim (5.2) holds for $k=1$. Then we assume that the claim (5.2) holds for some $k\in\mathbb{N}$. Note that $B^{k+1}=B\cdot B^{k}$. By (3.10) we have
  • for $1\leq j\leq m$, $b_{1,j}^{(k+1)}=bb_{1,j}^{(k)}+b_{2,j}^{(k)}=bv_{k+1-j}-v_{k+2-j-m}=v_{(k+1)-j+1}$, where the last equality follows from (ii);
  • for $1<i<m$ and $1\leq j\leq m$, $b_{i,j}^{(k+1)}=b_{i+1,j}^{(k)}=-v_{k+1+i-j-m}$;
  • for $1\leq j\leq m$, $b_{m,j}^{(k+1)}=-b_{1,j}^{(k)}=-v_{k+1-j}$.
This means that the claim (5.2) also holds for $k+1$. By induction on $k$, we prove the claim (5.2).
For $k\in\mathbb{N}$, write $C_{k}=AB^{k}=(c_{i,j}^{(k)})_{1\leq i,j\leq m}$. Then by (3.10) we have
(5.3) $\begin{cases}c_{i,j}^{(k)}=(b-1)v_{k-j+1}-v_{k-j+1-(m-i)}\;\text{for}\;1\leq i<m,1\leq j\leq m,\\ c_{m,j}^{(k)}=(b-1)v_{k-j+1}\;\text{for}\;1\leq j\leq m.\end{cases}$
By (ii) and (5.3), we first have that the matrix $AB^{k}$ is nonnegative for all $k\in\mathbb{N}$. Thus, the matrix $A^{p}B^{q}$ is nonnegative for all $p\in\mathbb{N},q\in\mathbb{N}$. For $k\geq m-1$, by (i) and (5.3) we obtain that all entries of the matrix $AB^{k}$ are positive. Thus, the matrix $A^{p}B^{q}$ is positive for $p\in\mathbb{N},q\in\mathbb{N}_{\geq m-1}$.
Next, fix $1\leq k<m-1$. Note that $v_{-m}=-1$. By (i) and (5.3) we have
$c_{i,j}^{(k)}>0\quad\text{for}\;j\leq k+1\;\text{or}\;j=k+1+i.$
Using the fact that $c_{i,1}^{(k)}>0$ for $1\leq i\leq m$, $c_{1,j}^{(k)}>0$ for $1\leq j\leq k+1$, and $c_{i,k+1+i}^{(k)}>0$ for $1\leq i\leq m-1-k$, one can check that the matrix $AB^{k}$ is primitive. For $\ell\in\mathbb{N}$, write $A^{\ell}B^{k}=(u_{i,j}^{(\ell)})_{1\leq i,j\leq m}$. By induction on $\ell$, we can show that $u_{i,1}^{(\ell)}>0$ for $1\leq i\leq m$, $u_{1,j}^{(\ell)}>0$ for $1\leq j\leq\min\{k+\ell,m\}$, and $u_{i,k+\ell+i}^{(\ell)}>0$ for $1\leq i\leq\max\{1,m-\ell-k\}$. This implies that the matrix $A^{\ell}B^{k}$ is primitive for all $\ell\in\mathbb{N}$.
Therefore, for any $p,q\in\mathbb{N}$ the matrix $A^{p}B^{q}$ is a primitive nonnegative matrix. ∎
Proof of Theorem 1.8.
Fix $p,q\in\mathbb{N}$. By Lemma 5.2, the matrix $A^{p}B^{q}$ is a primitive nonnegative matrix. Let $\lambda_{p,q}$ be the largest eigenvalue of the matrix $A^{p}B^{q}$. For $\omega\in L_{p,q}$, by Lemma 3.2 we have
$\mathbf{N}^{\omega}_{k(p+q)}=\mathbf{N}^{\omega}_{0}(A^{p}B^{q})^{k}\quad\forall k\in\mathbb{N}.$
As we have done in Lemma 3.4, there exists a constant $C>1$ such that
$C^{-1}\lambda_{p,q}^{k}\leq\|\mathbf{N}^{\omega}_{k(p+q)}\|\leq C\lambda_{p,q}^{k}\quad\forall k\in\mathbb{N}.$
By Proposition 3.1, the set $K^{\omega}$ is $\frac{\log\lambda_{p,q}}{(p+q)\log b}$-Ahlfors regular.
Next, we consider $\lambda_{np,nq}$ for $n\in\mathbb{N}$. Note that for $\omega\in L_{np,nq}$, we have $\mathbf{N}^{\omega}_{kn(p+q)}=\mathbf{N}^{\omega}_{0}(A^{np}B^{nq})^{k}$ for $k\in\mathbb{N}$. By Lemma 3.4, we have
$C_{1}^{-2k}\lambda^{knp}\eta^{knq}\leq\frac{\|\mathbf{N}^{\omega}_{kn(p+q)}\|}{\|\mathbf{N}^{\omega}_{0}\|}\leq C_{1}^{2k}\lambda^{knp}\eta^{knq}.$
On the other hand,
$\lim_{k\to\infty}\frac{\log\|\mathbf{N}^{\omega}_{kn(p+q)}\|}{kn(p+q)}=\frac{\log\lambda_{np,nq}}{n(p+q)}.$
Thus, we obtain
$\frac{p\log\lambda}{p+q}+\frac{q\log\eta}{p+q}-\frac{2\log C_{1}}{n(p+q)}\leq\frac{\log\lambda_{np,nq}}{n(p+q)}\leq\frac{p\log\lambda}{p+q}+\frac{q\log\eta}{p+q}+\frac{2\log C_{1}}{n(p+q)}.$
It follows that
$\lim_{n\to\infty}\frac{\log\lambda_{np,nq}}{n(p+q)}=\frac{p\log\lambda}{p+q}+\frac{q\log\eta}{p+q}.$
This means that the set
$\bigg\{\frac{\log\lambda_{p,q}}{(p+q)\log b}:p,q\in\mathbb{N}\bigg\}$
is dense in $\log_{b}\eta,\log_{b}\lambda]$. ∎
At the end of this section, we give a new pattern which also results in Ahlfors regularity.
Example 5.3.
Let $m\in\mathbb{N}_{\geq 3}$ and $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$, where $\omega^{k}=\omega^{k}_{1}\omega^{k}_{2}\ldots\omega^{k}_{m}\in D_{b}^{m}$. Suppose that for $k\in\mathbb{N}_{\geq m-1}$, we have
(5.4) $\begin{cases}\omega^{k}_{1}\ldots\omega^{k}_{m-j}\neq\omega^{k-j}_{j+1}\ldots\omega^{k-j}_{m}\quad\forall 1\leq j\leq m-2,\\ \omega^{k}_{1}=\omega^{k-m+1}_{m}.\end{cases}$
Note first that
[…]
where the union is disjoint. So, we obtain
(5.5) $\begin{split}|\Sigma_{k+1}^{\omega}|-(b-1)|\Sigma_{k}^{\omega}|&=\#\left\{d_{1}\ldots d_{k}\in\Sigma_{k}^{\omega}:d_{1}\ldots d_{k}\omega_{m}^{k-m+1}\in\Sigma_{k+1}^{\omega}\right\}\\ &=\#\left\{d_{1}\ldots d_{k}\in\Sigma_{k}^{\omega}:d_{1}\ldots d_{k}\omega_{1}^{k}\in\Sigma_{k+1}^{\omega}\right\},\end{split}$
where the last equality follows from $\omega^{k}_{1}=\omega^{k-m+1}_{m}$. On the other hand,
$\Sigma_{k+m-1}^{\omega}\times D_{b}=\Sigma_{k+m}^{\omega}\cup\big\{d_{1}\ldots d_{k}\omega_{1}^{k}\ldots\omega_{m}^{k}:d_{1}\ldots d_{k}\omega_{1}^{k}\ldots\omega_{m-1}^{k}\in\Sigma_{k+m-1}^{\omega}\big\},$
where the union is disjoint. Thus, we have
(5.6) $b|\Sigma_{k+m-1}^{\omega}|-|\Sigma_{k+m}^{\omega}|=\#\big\{d_{1}\ldots d_{k}\omega_{1}^{k}\ldots\omega_{m}^{k}:d_{1}\ldots d_{k}\omega_{1}^{k}\ldots\omega_{m-1}^{k}\in\Sigma_{k+m-1}^{\omega}\big\}.$
For any $d_{1}\ldots d_{k}\in\Sigma_{k}^{\omega}$, by (5.4) we have
$d_{1}\ldots d_{k}\omega_{1}^{k}\in\Sigma_{k+1}^{\omega}\;\;\text{if and only if}\;\;d_{1}\ldots d_{k}\omega_{1}^{k}\ldots\omega_{m-1}^{k}\in\Sigma_{k+m-1}^{\omega}.$
Therefore, by (5.5) and (5.6) we obtain $|\Sigma_{k+1}^{\omega}|-(b-1)|\Sigma_{k}^{\omega}|=b|\Sigma_{k+m-1}^{\omega}|-|\Sigma_{k+m}^{\omega}|$, i.e.,
$|\Sigma_{k+m}^{\omega}|=b|\Sigma_{k+m-1}^{\omega}|-|\Sigma_{k+1}^{\omega}|+(b-1)|\Sigma_{k}^{\omega}|\quad\forall k\geq m-1.$
Let $\gamma\in(b-1,b)$ be the root of equation $x^{m}-bx^{m-1}+x-(b-1)=0$. As we have done in Lemmas 3.3, 3.4 and Proposition 3.5, we can show that $\gamma$ is a Pisot number, and there exists a constant $C>1$ such that
$C^{-1}\gamma^{k}\leq|\Sigma_{k}^{\omega}|\leq C\gamma^{k}\quad\forall k\in\mathbb{N}.$
By Proposition 3.1, the set $K^{\omega}$ is $\frac{\log\gamma}{\log b}$-Ahlfors regular. Moreover, we have $\eta<\gamma<\lambda$.
Theorem 1.8 shows that for a countable and dense set of $s\in[\frac{\log\eta}{\log b},\frac{\log\lambda}{\log b}]$ there are infinitely many $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ such that $K^{\omega}$ is $s$-Ahlfors regular. We conjecture that this Ahlfors regular property should hold for all $s\in[\frac{\log\eta}{\log b},\frac{\log\lambda}{\log b}]$.
6. Applications in badly approximable numbers, non-recurrence points and joint spectral radius
In this section we will prove Theorems 1.9 and 1.11 for the set of badly approximable numbers and non-recurrence points, respectively, and prove Theorem 1.13 that the finiteness property holds for the joint spectral radius of $\left\{A_{\mathbf{d}}:\mathbf{d}\in D_{b}^{m}\right\}$.
Lemma 6.1.
Suppose that $m\in\mathbb{N}_{\geq 10}$. For any $\omega=\omega^{0}\omega^{1}\ldots,\tau=\tau^{0}\tau^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ we have
@@ 6. Applications in badly approximable numbers, non-recurrence points and joint spectral radius
(6.1) $\mathbf{d}^{(j)}\notin\bigcup_{i=0}^{q-1}\left\{\omega^{jq+i}_{q+1-i}\omega^{jq+i}_{q+2-i}\ldots\omega^{jq+i}_{q+\ell-i},\penalty 10000\ \tau^{jq+i}_{q+1-i}\tau^{jq+i}_{q+2-i}\ldots\tau^{jq+i}_{q+\ell-i}\right\}.$
@@
$K_{n}:=\left\{x\in[0,1):T_{b}^{k}(x)\notin B_{k}\quad\forall k\in\mathbb{N}_{\geq n}\right\}.$
@@
$\widetilde{K}_{n}:=T_{b}^{n}(K_{n})=\left\{y\in[0,1):T_{b}^{k}(y)\notin B_{k+n}\quad\forall k\in\mathbb{N}_{0}\right\}.$
@@
First we prove (i). Write $r_{n}=\mathrm{diam}\penalty 10000\ (B_{n})$ for $n\in\mathbb{N}_{0}$. Suppose that $r_{n}\to 0$ as $n\to\infty$. Then for any $m\in\mathbb{N}_{\geq 10}$ we can find $n_{0}\in\mathbb{N}$ such that $r_{n_{0}+k}<b^{-m}$ for all $k\in\mathbb{N}_{0}$. Take $n\geq n_{0}$. Then for any $k\in\mathbb{N}_{0}$ we can find two words $\mathbf{c}_{k},\mathbf{d}_{k}\in D_{b}^{m}$ such that $B_{k+n}\subset I_{\mathbf{c}_{k}}\cup I_{\mathbf{d}_{k}}$. It follows that
(6.3) $\widetilde{K}^{\omega,\tau}:=\left\{x\in[0,1):T_{b}^{k}(x)\not\in I_{\mathbf{c}_{k}}\cup I_{\mathbf{d}_{k}}\quad\forall k\in\mathbb{N}_{0}\right\}\subset\widetilde{K}_{n},$
where $\omega=\mathbf{c}_{0}\mathbf{c}_{1}\mathbf{c}_{2}\ldots,\tau=\mathbf{d}_{0}\mathbf{d}_{1}\mathbf{d}_{2}\ldots\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}$. Note that $\widetilde{K}^{\omega,\tau}$ and $K^{\omega,\tau}$ differ up to a countable set. By (6.3) and Lemma 6.1, we obtain that
@@
Next we prove (ii). Suppose $\liminf_{n\to\infty}r_{n}=r>0$. Choose $m\in\mathbb{N}$ such that $2b^{-m}<r$. Then there exists $n_{1}\in\mathbb{N}$ such that $r_{n_{1}+k}>2b^{-m}$ for all $k\in\mathbb{N}_{0}$. Take $n\geq n_{1}$. For each $k\in\mathbb{N}_{0}$ we can choose $\mathbf{d}_{k}\in D_{b}^{m}$ such that $I_{\mathbf{d}_{k}}\subset B_{k+n}$. Then we have
(6.4) $\widetilde{K}_{n}\subset\left\{x\in[0,1):T_{b}^{k}(x)\not\in\mathrm{int}\big(I_{\mathbf{d}_{k}}\big)\quad\forall k\in\mathbb{N}_{0}\right\}=:\widetilde{K}^{\omega},$
where $\omega=\mathbf{d}_{0}\mathbf{d}_{1}\mathbf{d}_{2}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$. Note that the sets $\widetilde{K}^{\omega}$ and $K^{\omega}$ differ up to a countable set. So, by Theorem 1.6 we obtain
@@
$E_{n}(\phi):=\left\{x\in[0,1):|T_{b}^{k}(x)-x|\geq\phi(k)\quad\forall k\geq n\right\}.$
@@
(6.5) $\dim_{H}E(\phi)=\sup_{n\geq 1}\dim_{H}E_{n}(\phi)=\lim_{n\to\infty}\dim_{H}E_{n}(\phi).$
First we prove (i). Suppose $\phi(n)\to 0$ as $n\to\infty$. For any $m\in\mathbb{N}_{\geq 10}$ we can find $n_{0}\in\mathbb{N}$ such that $\phi(n)<b^{-m}$ for all $n\geq n_{0}$. For $n\geq\max\{m,n_{0}\}$ we have
$E_{n}(\phi)\cap[0,b^{-m}]\supset\left\{x\in[0,b^{-m}]:T_{b}^{k}(x)\not\in[0,2b^{-m})\quad\forall k\geq n\right\}\supset\widetilde{E}_{n},$
where
By Lemma 6.1 we obtain that
$\dim_{H}E_{n}(\phi)\geq\dim_{H}E_{n}(\phi)\cap[0,b^{-m}]\geq\dim_{H}\widetilde{E}_{n}\geq\dim_{H}(T_{b}^{n}(\widetilde{E}_{n}))\geq 1-\frac{2}{m-2}\Big(\frac{\log m}{\log b}+1\Big).$
It follows from (6.5) that
Letting $m\to\infty$ we conclude that $\dim_{H}E\left(\phi\right)=1$ as desired.
Next we prove (ii). Suppose on the contrary that $\liminf_{n\to\infty}\phi(n)>0$. Note that $\phi(n)>0$ for all $n\in\mathbb{N}$. Then $r:=\displaystyle\inf_{n\geq 1}\phi(n)>0$. Choose $m\in\mathbb{N}_{\geq 10}$ such that $b^{-m}<r$. Then for any word $\mathbf{d}=d_{1}d_{2}\ldots d_{m}\in D_{b}^{m}$, we have for $n\in\mathbb{N}$,
$E_{n}(\phi)\cap I_{\mathbf{d}}\subset\left\{x\in[0,1):T_{b}^{k}(x)\notin I_{\mathbf{d}}\quad\forall k\geq n\right\}\subset T_{b}^{-n}(\widetilde{K}^{\omega})\quad\textrm{with}\quad\omega=\mathbf{d}^{\infty}\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}},$
@@
On the other hand, suppose that $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots=(\omega^{0}\omega^{1}\ldots\omega^{n-1})^{\infty}\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}$ is a periodic and progressively overlapping sequence. By Lemma 2.3 and Proposition 3.5, we have